# Hom-orthogonal modules and brick-Brauer-Thrall conjectures

### v1

@@ Abstract.
For finite dimensional algebras over algebraically closed fields, we study the sets of pairwise Hom-orthogonal modules and obtain new results on some open conjectures on the behaviour of bricks and several related problems, which we generally refer to as brick-Brauer-Thrall (bBT) conjectures. Using some algebraic and geometric tools, and in terms of the notion of Hom-orthogonality, we find necessary and sufficient conditions for the existence of infinite families of bricks of the same dimension. This sheds new lights on the bBT conjectures and we prove many of them for some new families of algebras. Our results imply some interesting algebraic and geometric characterizations of brick-finite algebras as conceptual generalizations of local algebras. We also verify the bBT conjectures for any algebra whose Auslander-Reiten quiver has a generalized standard component. On the one hand, our work relates and develops some recent studies of the bBT conjectures which are conducted in several directions. On the other hand, it extends some earlier results of Chindris-Kinser-Weyman, where the algebras with preprojective components were treated.
@@ 1. Introduction and Main Results
Theorem A: For algebra $A$, if $\mathcal{S}=\{X_{i}\}_{i\in I}$ is a Hom-orthogonal set in $\operatorname*{mod}A$, then there exists a Hom-orthogonal set of bricks $\mathcal{S}_{b}=\{B_{i}\}_{i\in I}$, where each $B_{i}$ is a submodule and a quotient module of $X_{i}$. Consequently, we have
@@
    If $A$ is a brick-finite algebra of rank $n$, there is no Hom-orthogonal set of modules with $n+1$ elements. Moreover, there exists a unique semibrick of size $n$, given by the simple modules in $\operatorname*{mod}A$.
@@
The next theorem summarizes some of our results on Hom-orthogonality in a more geometric setting. More precisely, for each representation variety $\operatorname*{rep}(A,\underline{d})$, we study the interaction between the orbits of an irreducible component $\mathcal{Z}$ of $\operatorname*{rep}(A,\underline{d})$. The undefined notations and terminology are recalled in Sections 2 and 4.
Theorem B: Let $A$ be an algebra and $\mathcal{Z}$ an irreducible component of $A$. If $\mathcal{Z}$ has a dense orbit, then for each $X$ and $Y$ in $\mathcal{Z}$, we have $\operatorname{Hom}_{A}(X,Y)\neq 0$. Furthermore, the following are equivalent:
@@
Note that $A$ is a local algebra if and only if, for any pair of non-zero modules $X$ and $Y$, we have $\operatorname{Hom}_{A}(X,Y)\neq 0$. The last assertion of Theorem B implies that brick-finite algebras can be seen as a conceptual generalization of local algebras (see Corollary 4.3). This further allows us to rephrase the Second brick-Brauer-Thrall conjecture (Conjecture 2.1) as follows: $A$ is brick-finite if and only if for each brick component $\mathcal{Z}$, and every pair $X$ and $Y$ in $\mathcal{Z}$, we have $\operatorname{Hom}_{A}(X,Y)\neq 0$.
Our last theorem is concerned with some interesting connections between the generalized standard components in the Auslander-Reiten quivers and the infinite families of bricks of the same dimension. In particular, for those algebras which admit a generalized standard component, the following theorem verifies several open brick-Brauer-Thrall conjectures listed in Section 2.2. The necessary materials and terminology are recalled in Sections 2 and 5.
@@
The stable tubes in the Auslander-Reiten quiver $\Gamma_{A}$ are known to be decisive in the representation theory of $A$. For instance, the Second Brauer-Thrall conjecture (now theorem), together with some fundamental results of Crawley-Boevey [C-B], imply that a tame algebra $A$ admits a $1$-parameter family of indecomposables if and only if $\Gamma_{A}$ has a stable tube. This is the case, if and only if $\Gamma_{A}$ contains an infinite family of homogeneous tubes whose mouths consist of modules of the same dimension. As a consequence of the previous theorem, we obtain a brick-analogue of these important results over tame algebras.
@@
As explained in the following sections, the Second brick-Brauer-Thrall ($2nd$ bBT) conjecture has played a central role in our recent studies. Although this problems can be treated in more technical settings (for instance, see [Mo1, Mo2], [MP1, MP2, MP3], [STV], [Pf1, Pf2]), we believe the novelty of the $2nd$ bBT conjecture is more apparent from its non-technical articulation: If $A$ is brick-infinite, then there exists an infinite family of bricks of the same dimension. In fact, some of the other open conjectures listed in Section 2.2 either follow from the $2nd$ bBT (see Remark 3.5), or are shown to be equivalent to it in certain settings (see [MP2] and [Pf2]). Let us conclude this part by pointing out that the basic notion of Hom-orthogonality that we adopt in this work leads to some new equivalent conditions which should provide fresh impetus to the study of several challenging problems. In particular, for an algebra $A$, our results imply that the following are equivalent:
@@
In the above set of equivalences, if one (and thus all) of these conditions holds, we can choose $d=d^{\prime}$ such that all the statements hold for $d$. Because these implications are consequences of our main results, we do not provide a separate proof. Instead, we observe that $(1)\rightarrow(2)$ follows from Proposition 3.4, and $(2)\rightarrow(3)$ follows from some standard geometric observations, as in Corollary 4.4, and $(3)\rightarrow(4)\rightarrow(1)$ follow from Theorem 4.2.
We finish this section with an outline of the content of the following sections. In Section 2 we recall some background materials and the open conjectures related to the scope of our studies, which we generally call the brick-Brauer-Thrall (bBT) conjectures. In Section 3, we consider the notion of Hom-orthogonality in the category of finite dimensional modules and prove Theorem A. In Section 4, we treat the notion of orthogonality in the irreducible components of representation varieties and prove Theorem B. In Section 5, we consider the connection between the components of the Auslander-Reiten components and the behavior of bricks, and prove Theorem C and Corollary D.
@@ 2.1. Representation varieties and their components
Let $\underline{d}$ be a non-zero dimension vector and $\mathcal{Z}$ in $\operatorname*{Irr}(A,\underline{d})$. For a pair of representations $X$ and $Y$ in $\mathcal{Z}$, we say the orbits $O_{X}$ and $O_{Y}$ are orthogonal provided that $X$ and $Y$ are Hom-orthogonal modules, that is $\operatorname{Hom}_{A}(X,Y)=\operatorname{Hom}_{A}(Y,X)=0$. More generally, for $\mathcal{Z}\in\operatorname*{Irr}(A)$, we say $\mathcal{Z}$ has a family of orthogonal orbits if $\mathcal{Z}$ contains a family of pairwise orthogonal orbits as above. Let us also recall that if $M$ and $N$ are two modules that belong to $\mathcal{Z}$, we say $M$ degenerates to $N$, or equivalently $N$ is called a degeneration of $M$, provided that $N\not\simeq M$ and $N$ belongs to $\overline{O_{M}}$, that is, $O_{M}$ and $O_{N}$ are distinct orbits in $\mathcal{Z}$ and $O_{N}\subsetneq\overline{O_{M}}$. The following theorem gives an elegant algebraic characterization of this geometric notion.
Theorem 2.2 ([Zw])
Let $\mathcal{Z}$ be an irreducible component of $A$. For any pair of modules $M$ and $N$ in $\mathcal{Z}$, the following are equivalent:
  1. (1)
    $M$ degenerates to $N$;
  2. (2)
    there exists a short exact sequence $0\rightarrow N\rightarrow M\oplus Z\rightarrow Z\rightarrow 0$ in $\operatorname*{mod}A$, for some $Z$;
  3. (3)
    there exists a short exact sequence $0\rightarrow Z\rightarrow M\oplus Z\rightarrow N\rightarrow 0$ in $\operatorname*{mod}A$, for some $Z$.
@@
The next conjecture is originally inspired by the geometry of the fan induced by $\tau$-tilting theory. As described in [DIJ], to an algebra $A$, one can associate a polyhedral fan constructed by the $g$-vectors of the indecomposable $\tau$-rigid modules and the shifts of the indecomposable projective modules, which we simply call the $\tau$-tilting fan of $A$, and denote it by $\mathcal{F}_{A}$. This is a conceptual analogue of the $g$-vector fan studied in cluster algebras. If $A$ is of rank $n$, then it is shown that $A$ is brick-finite if and only if $\mathcal{F}_{A}$ is a complete fan in $\mathbb{R}^{n}$, namely, $\mathcal{F}_{A}=\mathbb{R}^{n}$. Recently there has been an extensive study of those algebras which are not brick-finite but their $\tau$-tilting fan manifest interesting properties (see [PY], [AI], and the references therein). The following statement was first posed as a question in the work of Demonet [De] and it is open in full generality. There has been some recent work on the connections between the $2nd$ bBT conjecture and the following conjecture (see [MP2] and [Pf2]).
@@
The next conjecture that we mention here predates the above conjectures, but closely relates to our studies of bricks. The following conjecture is originally phrased in a slightly more technical language, and in terms of dimension of the algebras of semi-invariants on irreducible components in $\operatorname*{Irr}(A)$ (see [CKW, Conjecture 5.4]). However, for simplicity, below we present it in the language of stable modules. Recall that, for $\theta\in K_{0}(\operatorname*{proj}A)\simeq\mathbb{Z}^{n}$, a module $M$ in $\operatorname*{mod}A$ is $\theta$-stable if $\theta(\underline{\dim}M)=0$, and moreover $\theta(\underline{\dim}N)<0$, for each nonzero submodule $N\subsetneq M$.
@@
If $A$ is brick-infinite, there exists some $\theta\in K_{0}(\operatorname*{proj}A)$, such that there are infinitely many $\theta$-stable modules of the same dimension.
@@ 2.3. Generalized standard components
Starting from an algebra $B$ and a family of infinite generalized standard components in $\Gamma_{B}$, as in [MS1] one can iteratively apply a sequence of 10 admissible operations to get new translation quivers. For each admissible operation on the translation quiver, there is a corresponding admissible operation on the algebra under consideration (for these operations on translation quivers and algebras, see [MS1, Section 2]). A connected translation quiver $\mathcal{C}$ is called a generalized multicoil if it can be obtained from a finite family $\{\mathcal{T}_{1},\cdots,\mathcal{T}_{r}\}$ of stable tubes via an iterated application of the admissible translation operations. We remark that for any generalized multicoil $\mathcal{C}$, the cyclic part of $\mathcal{C}$ is infinite, connected and cofinite in $\mathcal{C}$. Thus, any generalized multicoil is a connected almost cyclic translation quiver.
@@
As shown in [MS3, Theorem 1.2], for an infinite generalized standard component $\mathcal{C}$ in $\Gamma_{A}$, one can consider three quotient algebras of $A$, denoted by $A^{(lt)}_{\mathcal{C}}$, $A^{(rt)}_{\mathcal{C}}$, and $A^{(c)}_{\mathcal{C}}$, respectively called the left tilted algebra of $\mathcal{C}$, the right tilted algebra of $\mathcal{C}$, and the coherent algebra of $\mathcal{C}$. In particular, each one of $A^{(lt)}_{\mathcal{C}}$ and $A^{(rt)}_{\mathcal{C}}$ is a finite product of some tilted algebras. Moreover, $A^{(c)}_{\mathcal{C}}=A^{(c)}_{1}\times\cdots\times A^{(c)}_{p}$, where, for each $1\leq i\leq p$, the algebra $A^{(c)}_{i}$ is a generalized multicoil enlargement of an algebra $B^{(c)}_{i}$ with a faithful family of pairwise orthogonal generalized standard stable tubes in $\Gamma_{B^{(c)}_{i}}$ (for details, see [MS3, Sections 2 &\&& 3]). As discussed in Section 5, this important result allows us to employ several reduction techniques in the study of bricks over those algebras which admit a generalized standard components.
@@ 3. Hom-orthogonal modules and brick-finiteness
It is expected that the converse of part $(1)$ of the above corollary is true in general (see Remark 3.5). Moreover, if the infinite Hom-orthogonal set in part $(2)$ has bounded dimensions, then the same holds for the infinite set of bricks, which would prove the $2nd$ bBT conjecture (Conjecture 2.1). Note that, if $A$ is brick-infinite, then by the $2nd$ BT conjecture (now theorem), there exists an ascending sequence of integers $0<d_{1}<d_{2}<\cdots$ such that every $\operatorname*{ind}(A,d_{i})$ is an infinite set. Thus, to prove the $2nd$ bBT conjecture for a brick-infinite algebra $A$, it suffices to show that one of such infinite families $\operatorname*{ind}(A,d_{i})$ contains an infinite Hom-orthogonal subset. In Section 4 we prove a stronger version of this fact.
Through some geometric arguments in representation varieties, in the next proposition we obtain a more refined version of some known results on brick components (compare [G+, Corollary 1.6]). In the following, for $X$ in $\operatorname*{mod}A$, we set ${X}^{\perp_{0}}:=\{Y\in\operatorname*{mod}A\,|\,\operatorname{Hom}_{A}(X,Y)=0\}$. We dually define $\,{}^{\perp_{0}}{X}$. The next result gives a particular attention to the Hom-orthogonal bricks of the same dimension and immediately implies part $(2)$ of Theorem A.
@@
By a systematic treatment of the $2nd$ bBT conjecture via $\tau$-tilting theory, in [MP3] we reduced this open conjecture to a certain family of algebras. Before we state the connections between our new results and this reduction, recall that we say $A$ is minimal brick-infinite algebra if $A$ is brick-infinite, but each proper quotient of $A$ is brick-finite. We note that if $B$ is a brick-infinite quotient of $A$ and satisfies one of the conjectures from Section 2.2, then that conjecture immediately holds for $A$. In particular, to prove the (stable) $2nd$ bBT conjecture for a brick-infinite algebra, it obviously suffices to verify that over one of its minimal brick-infinite quotients. Moreover, [MP3, Theorem 1.4] gives a non-trivial reduction on this open problem: To settle the $2nd$ bBT conjecture in full generality, it is sufficient to verify it over those minimal brick-infinite algebras for which almost all bricks are faithful. Such reductive arguments can be similarly employed in the study of the other brick-Brauer-Thrall conjectures. For instance, in [MP1], we explicitly described the minimal brick-infinite biserial algebras. From that, one can verify all the conjectures in Section 2.2 for arbitrary biserial algebras. On the other hand, we can strengthen our earlier results on some minimal brick-infinite algebras.
@@
We first recall that a wide subcategory $\mathcal{W}$ in $\operatorname*{mod}A$ is functorially finite if and only if $\mathcal{W}$ is equivalent to $\operatorname*{mod}B$, for a finite dimensional algebra $B$ (for example, see [En2, Proposition 4.12]). In particular, if $A$ is brick-finite, then each wide subcategory is of this form, that is, every wide subcategory $\mathcal{W}$ in $\operatorname*{mod}A$ is functorially finite (see [MS, Corollary 3.11]). Observe that if $A$ is brick-infinite, Conjecture 2.1 implies that there is an infinite family of bricks of the same dimension, say $\{X_{i}\}_{i\in\mathbb{Z}}$ in $\operatorname*{brick}(A)$. Moreover, Proposition 3.4 yields an infintie sembrick as a subset of $\{X_{i}\}_{i\in\mathbb{Z}}$. Now, if we consider the wide subcategory $\mathcal{W}$ generated by this infinite semibrick, then we obtain a wide subcategory $\mathcal{W}$ in $\operatorname*{mod}A$ which has infinitely many simple objects. Evidently, $\mathcal{W}$ cannot be equivalent to the module category of a finite dimensional algebra, that is, $\mathcal{W}$ is not functorially finite. ∎

### v2

@@ Abstract.
For finite dimensional algebras over algebraically closed fields, we study the sets of pairwise Hom-orthogonal modules and obtain new results on some open conjectures on the behaviour of bricks and several related problems, which we generally refer to as brick-Brauer-Thrall (bBT) conjectures. Using some algebraic and geometric tools, and in terms of the notion of Hom-orthogonality, we find necessary and sufficient conditions for the existence of infinite families of bricks of the same dimension. This sheds new light on the bBT conjectures and we prove some of them for new families of algebras. Our results imply some interesting algebraic and geometric characterizations of brick-finite algebras as conceptual generalizations of local algebras. We also verify the bBT conjectures for any algebra whose Auslander-Reiten quiver has a generalized standard component, which particularly extends some results of Chindris-Kinser-Weyman on the algebras with preprojective components.
@@ 1. Introduction and Main Results
Theorem A: For algebra $A$, if $\mathcal{S}=\{X_{i}\}_{i\in I}$ is a Hom-orthogonal set in $\operatorname*{mod}A$, then there exists a Hom-orthogonal set of bricks $\mathcal{S}_{b}=\{B_{i}\}_{i\in I}$, where each $B_{i}$ is a submodule and a quotient module of $X_{i}$. Moreover, we have
@@
    If $A$ is a brick-finite algebra of rank $n$, then there is no Hom-orthogonal set of modules with $n+1$ elements. In addition, there exists a unique semibrick of size $n$, and it is given by the simple modules in $\operatorname*{mod}A$.
@@
The next proposition summarizes some of our results on Hom-orthogonality in a more geometric setting. More precisely, for each representation variety $\operatorname*{rep}(A,\underline{d})$, we study the interaction between the orbits of an irreducible component $\mathcal{Z}$ of $\operatorname*{rep}(A,\underline{d})$. The undefined notations and terminology are recalled in Sections 2 and 4.
Proposition B: Let $A$ be an algebra and $\mathcal{Z}$ an irreducible component of $A$. If $\mathcal{Z}$ has a dense orbit, then for each $X$ and $Y$ in $\mathcal{Z}$, we have $\operatorname{Hom}_{A}(X,Y)\neq 0$. Furthermore, the following are equivalent:
@@
The first assertion is proved in Lemma 4.1, and the complete proof of Proposition B is given in Section 4, where we also make some observations on connections to some recent related work in the geometric setting. Meanwhile, let us note that $A$ is a local algebra if and only if, for any pair of non-zero modules $X$ and $Y$, we have $\operatorname{Hom}_{A}(X,Y)\neq 0$. Hence, the last assertion of Proposition B implies that brick-finite algebras can be seen as a conceptual generalization of local algebras (see Corollary 4.3). This further allows us to rephrase the Second brick-Brauer-Thrall conjecture (Conjecture 2.1) as follows: $A$ is brick-finite if and only if for each brick component $\mathcal{Z}$, and every pair $X$ and $Y$ in $\mathcal{Z}$, we have $\operatorname{Hom}_{A}(X,Y)\neq 0$.
Our second main result is concerned with some interesting connections between the generalized standard components in the Auslander-Reiten quivers and the infinite families of bricks of the same dimension. In particular, for those algebras which admit a generalized standard component, the following theorem verifies several open brick-Brauer-Thrall conjectures listed in Section 2.2. The necessary materials and terminology are recalled in Sections 2 and 5.
@@
The stable tubes in the Auslander-Reiten quiver $\Gamma_{A}$ are known to be decisive in the representation theory of $A$. For instance, the Second Brauer-Thrall conjecture (now theorem), together with some fundamental results of Crawley-Boevey [C-B], imply that a tame algebra $A$ admits a $1$-parameter family of indecomposables if and only if $\Gamma_{A}$ has a stable tube. This is the case, if and only if $\Gamma_{A}$ contains an infinite family of homogeneous tubes whose mouths consist of modules of the same dimension. As a consequence of the previous theorem, and using some classical results on tame algebras, we obtain a brick-analogue of these important results over tame algebras.
@@
As explained in the following sections, the Second brick-Brauer-Thrall ($2nd$ bBT) conjecture has played a central role in our recent studies. Although this problem can be treated in more technical settings (for instance, see [Mo1, Mo2], [MP1, MP2, MP3], [STV], [Pf1, Pf2]), we believe the novelty of the $2nd$ bBT conjecture is more apparent from its non-technical articulation: If $A$ is brick-infinite, then there exists an infinite family of bricks of the same dimension. In fact, some of the other open conjectures listed in Section 2.2 either follow from the $2nd$ bBT (see Remark 3.5), or are shown to be equivalent to it in certain settings (see [MP2] and [Pf2]).
Let us remark that the basic notion of Hom-orthogonality that we adopt in this work leads to some equivalent conditions that should provide fresh impetus to the study of several challenging problems. In particular, for an algebra $A$, our results imply that the following are equivalent:
@@
We observe that, after applying some geometric arguments, some of the above implications can also be concluded from [G+, Theorem 1.5]. We further note that, in the above set of equivalences, if one (and thus all) of these conditions holds, we can choose $d=d^{\prime}$ so that all of the statements hold for $d$. Because these implications are consequences of our main results, separate proofs are not provided. Instead, let us remark that $(1)\rightarrow(2)$ follows from Proposition 3.4, and $(2)\rightarrow(3)$ follows from a standard geometric observation, and $(3)\rightarrow(4)\rightarrow(1)$ follow from Theorem 4.2.
We finish this section with an outline of the content of the following sections. In Section 2 we recall some background materials and the open conjectures related to the scope of our studies, which we generally call the brick-Brauer-Thrall (bBT) conjectures. In Section 3, we consider the notion of Hom-orthogonality in the category of finite dimensional modules and prove Theorem A. In Section 4, we treat the notion of orthogonality in the irreducible components of representation varieties and prove Proposition B. In Section 5, we consider the connection between the components of the Auslander-Reiten quiver and the behavior of bricks, and prove Theorem C and Corollary D.
@@ 2.1. Representation varieties and their components
The next conjecture is originally inspired by the geometry of the fan induced by $\tau$-tilting theory. As described in [DIJ], to an algebra $A$, one can associate a polyhedral fan constructed by the $g$-vectors of the indecomposable $\tau$-rigid modules and the shifts of the indecomposable projective modules, which we simply call the $\tau$-tilting fan of $A$, and denote it by $\mathcal{F}_{A}$. This is a conceptual analogue of the $g$-vector fan studied in cluster algebras. If $A$ is of rank $n$, then it is shown that $A$ is brick-finite if and only if $\mathcal{F}_{A}$ is a complete fan in $\mathbb{R}^{n}$, namely, $\mathcal{F}_{A}=\mathbb{R}^{n}$. Recently there has been an extensive study of those algebras which are not brick-finite but their $\tau$-tilting fan manifest interesting properties (see [PY], [AI], and the references therein). The following statement, thanks to the characterization of $\tau$-reduced components by Plamondon in [Pl], is equivalent to a question that was first posed in the work of Demonet [De] and it is open in full generality. There has been some recent work on the connections between the $2nd$ bBT conjecture and the following conjecture (see [MP2] and [Pf2]).
@@
The next conjecture that we mention here predates the above conjectures, but closely relates to our studies of bricks. It was originally phrased in a slightly more technical language, and in terms of dimension of the algebras of semi-invariants on irreducible components in $\operatorname*{Irr}(A)$ (see [CKW, Conjecture 5.4]). However, for simplicity, below we present it in the language of stable modules. Recall that, for $\theta\in K_{0}(\operatorname*{proj}A)\simeq\mathbb{Z}^{n}$, a module $M$ in $\operatorname*{mod}A$ is $\theta$-stable if $\theta(\underline{\dim}M)=0$, and moreover $\theta(\underline{\dim}N)<0$, for each nonzero submodule $N\subsetneq M$.
@@
If $A$ is brick-infinite, then there exists some $\theta\in K_{0}(\operatorname*{proj}A)$ such that there are infinitely many $\theta$-stable modules of the same dimension.
@@ 2.3. Generalized standard components
Starting from an algebra $B$ and a family of infinite generalized standard components in $\Gamma_{B}$, as in [MS1] one can iteratively apply a sequence of 10 admissible operations to get new translation quivers. For each admissible operation on the translation quiver, there is a corresponding admissible operation on the algebra under consideration (for these operations on translation quivers and algebras, see [MS1, Section 2]). A connected translation quiver $\mathcal{C}$ is called a generalized multicoil if it can be obtained from a finite family $\{\mathcal{T}_{1},\ldots,\mathcal{T}_{r}\}$ of stable tubes via an iterated application of the admissible translation operations. We remark that for any generalized multicoil $\mathcal{C}$, the cyclic part of $\mathcal{C}$ is infinite, connected and cofinite in $\mathcal{C}$. Thus, any generalized multicoil is a connected almost cyclic translation quiver.
@@
As shown in [MS3, Theorem 1.2], for an infinite generalized standard component $\mathcal{C}$ in $\Gamma_{A}$, one can consider three quotient algebras of $A$, denoted by $A^{(lt)}_{\mathcal{C}}$, $A^{(rt)}_{\mathcal{C}}$, and $A^{(c)}_{\mathcal{C}}$, respectively called the left tilted algebra of $\mathcal{C}$, the right tilted algebra of $\mathcal{C}$, and the coherent algebra of $\mathcal{C}$. In particular, each one of $A^{(lt)}_{\mathcal{C}}$ and $A^{(rt)}_{\mathcal{C}}$ is a finite product of some tilted algebras. Moreover, $A^{(c)}_{\mathcal{C}}=A^{(c)}_{1}\times\cdots\times A^{(c)}_{p}$, where, for each $1\leq i\leq p$, the algebra $A^{(c)}_{i}$ is a generalized multicoil enlargement of an algebra $B^{(c)}_{i}$ with a faithful family of pairwise orthogonal generalized standard stable tubes in $\Gamma_{B^{(c)}_{i}}$ (for details, see [MS3, Sections 2 &\&& 3]). As discussed in Section 5, this important result allows us to employ several reduction techniques in the study of bricks over those algebras which admit a generalized standard component.
@@ 3. Hom-orthogonal modules and brick-finiteness
It is expected that the converse of part $(1)$ of the above corollary is true in general (see Remark 3.5). Moreover, if the infinite Hom-orthogonal set in part $(2)$ has bounded dimensions, then the same holds for the infinite set of bricks, which would yield the $2nd$ bBT conjecture (Conjecture 2.1). Note that, if $A$ is brick-infinite, then by the $2nd$ BT conjecture (now theorem), there exists an ascending sequence of integers $0<d_{1}<d_{2}<\cdots$ such that every $\operatorname*{ind}(A,d_{i})$ is an infinite set. Thus, to prove the $2nd$ bBT conjecture for a brick-infinite algebra $A$, it suffices to show that one of such infinite families $\operatorname*{ind}(A,d_{i})$ contains an infinite Hom-orthogonal subset. In Section 4 we prove a stronger version of this fact.
Through some geometric arguments in representation varieties, in the next proposition we obtain a more refined version of some known results on brick components (compare [G+, Corollary 1.6]). In that paper, the authors have shown that if $\mathcal{Z}$ is a brick component without a dense orbit, then there is an open set in $\mathcal{Z}\times\mathcal{Z}$ consisting of pairs of orthogonal bricks. We note that projecting such an open set down onto $\mathcal{Z}$ does not necessarily yield a set of Hom-orthogonal bricks. That being the case, to obtain the desired set of Hom-orthogonal bricks in $\mathcal{Z}$, some further observations are made in the proof of the following proposition. Below, for $X$ in $\operatorname*{mod}A$, we set ${X}^{\perp_{0}}:=\{Y\in\operatorname*{mod}A\,|\,\operatorname{Hom}_{A}(X,Y)=0\}$. We dually define $\,{}^{\perp_{0}}{X}$. Note that, together with Corollary 3.3, the following proposition immediately implies part $(2)$ of Theorem A.
@@
By a systematic treatment of the $2nd$ bBT conjecture via $\tau$-tilting theory, in [MP3] we reduced this conjecture to a certain family of algebras. Before we state the connections between our new results and this reduction, recall that we say $A$ is minimal brick-infinite algebra if $A$ is brick-infinite, but each proper quotient of $A$ is brick-finite. We note that if $B$ is a brick-infinite quotient of $A$ and satisfies one of the conjectures from Section 2.2, then that conjecture immediately holds for $A$. In particular, to prove the (stable) $2nd$ bBT conjecture for a brick-infinite algebra, it obviously suffices to verify that over one of its minimal brick-infinite quotients. Moreover, [MP3, Theorem 1.4] gives a non-trivial reduction on this open problem: To settle the $2nd$ bBT conjecture in full generality, it is sufficient to verify it over those minimal brick-infinite algebras for which almost all bricks are faithful. Such reductive arguments can be similarly employed in the study of the other brick-Brauer-Thrall conjectures. For instance, in [MP1], we explicitly described the minimal brick-infinite biserial algebras. From that, one can verify all the conjectures in Section 2.2 for arbitrary biserial algebras. On the other hand, we can strengthen our earlier results on some minimal brick-infinite algebras.
@@
We first recall that a wide subcategory $\mathcal{W}$ in $\operatorname*{mod}A$ is functorially finite if and only if $\mathcal{W}$ is equivalent to $\operatorname*{mod}B$, for a finite dimensional algebra $B$ (for example, see [En2, Proposition 4.12]). In particular, if $A$ is brick-finite, then each wide subcategory is of this form, that is, every wide subcategory $\mathcal{W}$ in $\operatorname*{mod}A$ is functorially finite (see [MS, Corollary 3.11]). Observe that if $A$ is brick-infinite, Conjecture 2.1 implies that there is an infinite family of bricks of the same dimension, say $\{X_{i}\}_{i\in\mathbb{Z}}$ in $\operatorname*{brick}(A)$. Moreover, Proposition 3.4 yields an infinite semibrick as a subset of $\{X_{i}\}_{i\in\mathbb{Z}}$. Now, if we consider the wide subcategory $\mathcal{W}$ generated by this infinite semibrick, then we obtain a wide subcategory $\mathcal{W}$ in $\operatorname*{mod}A$ which has infinitely many simple objects. Evidently, $\mathcal{W}$ cannot be equivalent to the module category of a finite dimensional algebra, that is, $\mathcal{W}$ is not functorially finite. ∎